# Laminator: Verifiable ML Property Cards using Hardware-assisted Attestations

### v1

@@ Abstract
Regulations increasingly call for various assurances from machine learning (ML) model providers about their training data, training process, and the behavior of resulting models during inference. For better transparency, companies (e.g., Huggingface and Google) have adopted model cards and datasheets which describe different properties of the training datasets and models. In the same vein, we introduce the notion of an inference card to describe the properties of a given inference (e.g., binding output to the model and its corresponding input). We collectively refer to these as ML property cards.
@@
A malicious model provider can include false information in ML property cards, raising a need for verifiable ML property cards. We show how to realized them using property attestation, technical mechanisms by which a prover (e.g., a model provider) can attest different ML properties during training and inference to a verifier (e.g., an auditor). However, prior attestation mechanisms based purely on cryptography are often narrowly focused (lacking versatility) and inefficient. There is a need to efficiently attest different types properties across the ML model training and inference pipeline.
@@
Recent developments make it possible to run and even train models inside hardware-assisted trusted execution environments (TEEs), which can provide highly efficient attestation. We propose Laminator, the first framework for verifiable ML property cards using hardware-assisted ML property attestations to efficiently furnish attestations for various ML properties for training and inference. It scales to multiple verifiers, and is independent of the model configuration.
I Introduction
@@
Machine learning (ML) models are increasingly being deployed for various high-stakes applications like medical diagnosis, bank loans, and autonomous vehicles [9]. This has raised concerns about the different risks to data privacy, fairness and robustness [31]. Further, there are several emerging regulations to ensure that the training process and model’s behaviour during inference are as expected [22, 63, 14, 10, 12, 38, 32, 23]. For instance, a model provider may be required to convince a verifier (e.g., a regulator or potential customer) that a model meets the expected level of accuracy, fairness, privacy or robustness without disclosing the model [22, 23, 32, 38, 12] or demonstrate to an auditor that distributional properties of training data comply with regulations [14, 10] without disclosing the training data itself. The ability to demonstrate regulatory compliance without compromising confidentiality of models or training data will enable the development of a marketplace for ML models.
@@
Several companies, including Google and Huggingface, have adopted model cards [43] and datasheets [18, 52] for better transparency of the properties of ML models or training datasets. We propse a similar notion, inference cards for describing the properties of a specific inference, e.g., binding the output of an inference, to the input and the model. Collectively, we refer to them as ML property cards which are short documents a containing information about ML properties such as intended use, risks and limitations, training procedure, datasets, and metrics for evaluation (e.g., accuracy, fairness, robustness) [43]. These ML property cards are generated by the model provider who trains and/or deploys the model. This requires trusting them to honestly provide the information in the ML property cards. While trusting reputable companies which open-source their models may be reasonable, we need to provide mechanisms to attest information in the ML property cards for regulatory compliance. Further, as seen on Huggingface, trained models and their corresponding property cards from potentially malicious entities are also available in model marketplaces [44]. Hence, there is a need for verifiable ML property cards to ensure that the information contained in them is neither false nor tampered with.
@@
Huggingface provides a utility by which model providers can send their model to an endorser like Huggingface in order to compute some metrics, and create a verifiable property card that contains a signature to convince the verifier to ensure that a metric was verified by the endorser. However, this is not versatile, i.e., it is limited only to metrics, models and datasets implemented or allowed by the endorser. Hence, it is not possible to prove i) properties of sensitive datasets that cannot be disclosed to the endorser (e.g., demographic distribution, or certifying provenance of training dataset) ii) that a model was trained using a specific optimization metric (e.g., differential privacy [4], or group fairness [21]), or iii) that an output is the result of the model applied on a corresponding input during inference. We propose using ML property attestation, first introduced by Duddu et al. [11], which are technical mechanisms by which a prover (e.g., a model provider) can demonstrate relevant properties (including training-time and inference-time) of the model to a verifier (e.g., a client querying a model deployed as a service, a customer purchasing a model, or a regulatory authority).
@@
Current ML property attestation schemes have various drawbacks in terms of accuracy [11], efficiency [11, 17], or scalability to multiple verifiers [11]. Further, prior works are limited to specific attestations such as proof of training (that a model was trained on some dataset) [17, 5, 57, 62], distributional property attestations (showing the training dataset of a model satisfies distributional properties) [11, 8], or showing that a model was trained with some optimization [56, 34, 50, 55, 68]. There is a need to efficiently attest various different types of properties during ML training or inference. We identify efficiency, scalability, and versatility as important criteria for attestation schemes.
@@
Trusted Execution Environments (TEEs) already have the notion of hardware-assisted remote attestation to prove local system or software configuration to a remote verifier. The trusted computing research community has extended this to the notion of property attestation [54]. Recent developments by hardware vendors [26, 49], make it possible to run and train ML models efficiently inside TEEs. Hence, we identify TEEs as the choice for realizing efficient, scalable and versatile ML property attestation. We are the first to propose a software framework for hardware-assisted ML property attestation. We use TEEs to establish bindings between key properties of training dataset and the model during inference and training. These bindings are signed by the TEE’s secret attestation key, generating ML property attestations. When combined with signed external certificates from trusted entities, these attestations enable verifiers to draw conclusions about the model and training dataset properties during inference. These attestations can be then be used for verifiable ML property cards. We claim the following contributions: we
  1. 1.
@@
    propose inference cards which include properties specific to each inference (e.g., binding input-model-output) and the collective of ML property cards to include model cards, datasheets, and inference cards (Section IV)
  2. 2.
@@
    propose Laminator  [^1] We will make the code publicly available after publication., the first framework to generate verifiable ML property cards using hardware-assisted attestations throughout training and inference (Section V)
  3. 3.
    design, implement and evaluate four different types of attestations on Intel SGX as part of Laminator (Section VI)
II Background
@@
We present the ML notations used in this work (Section II-A), details about model and dataset cards (Section II-B), and trusted execution environments (Section II-C).
II-A Machine Learning Notations
@@
Consider an ML model architecture configuration $\mathcal{M}_{ar}$ indicating the number and types of layers, and a training configuration file $\mathcal{T}$ which includes the objective function, learning rate, weight decay, and other hyperparameters such as number of epochs. Additionally, consider a dataset $\mathcal{D}_{tr}\sim\mathbb{D}_{\textsc{p}}$ where $\mathbb{D}_{\textsc{p}}$ is an underlying distribution satisfying some distributional properties p (for example, the ratio of females in the distribution). Since, we use i.i.d sampling to get $\mathcal{D}_{tr}$, it satisfies p. We obtain model $\mathcal{M}$ by training as per the description in $\mathcal{T}$ using $\mathcal{M}_{ar}$ on $\mathcal{D}_{tr}$, i.e., $\mathcal{M}\leftarrow\text{Train}(\mathcal{D}_{tr},\mathcal{M}_{ar},\mathcal{T})$. During inference, $\mathcal{M}$ takes an input $\mathcal{I}$ and output $\mathcal{O}$, i.e., $\mathcal{O}$leftarrow $\mathcal{M}$($\mathcal{I}$). Additionally, $\mathcal{M}$ is evaluated with respect to accuracy, fairness or robustness on some unseen test dataset $\mathcal{D}_{te}$ sampled from same distribution as $\mathcal{D}_{tr}$. We assume that both $\mathcal{D}_{tr}$ and $\mathcal{D}_{te}$ are composed of data records $\mathcal{I}$, the corresponding classification label $y$, and some sensitive attribute $z$. We summarize different notations used in the rest of the paper in Table I.
TABLE I: Summary of notations and their descriptions.
@@
II-B Model Cards and Datasheets
@@
Model cards are short documents containing additional information about different static model properties [43]. These include properties prior to deployment of the model as a service to clients to query via an API reference. They were designed to provide better transparency into the design of ML model which includes training and evaluation of the model prior to deployment for inference. These have been extensively adopted by different companies such as Google [20] and Huggingface [25]. Model cards contain a variety of additional information about ML properties: i) general propertiesincluding information about the model’s intended use, risks such as bias, and other limitations ii) training propertiesincluding training datasets used, training hyperparameters, and model architecture iii) evaluation propertiesincluding information about the different test datasets and metrics used for measuring accuracy, precision, recall, fairness, and robustness.
Datasheets describe different properties corresponding to the training dataset  [18, 52]. Datasheets, also known as dataset cards, have been used extensively by companies such as Huggingface [1]. They describe collection process and sources from which the raw data was collected, pre-processing including cleaning and labelling to obtain training dataset from raw data, intended use, composition of dataset, and demographic subpopulations and their distribution.
@@
II-C Trusted Execution Environments (TEEs)
@@
TEEs [36] provide an environment in which security-critical code (trusted applications, or TAs) can run without outside interference from other TAs, or the rest of the system (the Rich Execution Environment, or REE, which contains the operating system and normal software). TEEs can be described in terms of three main properties: isolation, secure storage, and remote attestation [36]. Isolation refers to the inability for outside software to interfere with the operation of a TA, or to read its confidential state. Secure storage refers to persistent storage made available to TAs that resists interference by other software. Remote attestation refers to the ability of a TA to prove some aspect of a TEE’s state to a remote party, e.g., that the secret key of a particular keypair is held by a TA whose code has a particular cryptographic hash.
In this work, we will focus on the use of Intel’s Software Guard Extensions (SGX) [2] to enable TEEs, but the work equally applies to other TEE technologies that provide both isolation and remote attestation. SGX provides enclave-type TEEs, meaning that individual applications can instantiate a TEE (an enclave) in their own memory space, running a TA of their choice. SGX prevents the host application, the operating system, or any other code outside the enclave, from reading or writing to the enclave’s memory directly. Communication with the enclave is via well-defined entry and exit points, as well as by the TA reading and writing the host application’s memory outside of the enclave.
SGX enclaves can remotely attest themselves by obtaining a quote, which functions as a signature over TA-provided ‘user data’, along with a certification of the enclave identity that produced it. This enclave identity includes MRENCLAVE, a hash of the TA’s initial memory image, MRSIGNER, a hash of the public key that signed the TA, whether the enclave is in debug mode, versioning information, and various other data [2, EREPORT].
@@
A verifier can evaluate the validity of the quote, and use its knowledge of the code whose hash is contained in the quote’s MRENCLAVE (and specifically, the circumstances under which it will produce a quote) to reliably determine properties of the TA’s state at the time that it produced the quote. In particular, if a particular TA produces quotes only over the result of a particular computation, then the verifier can be assured that such a computation was performed, and yielded the result contained in the quote. This can be used to prove arbitrary properties of data available to and computations performed by a TA.
II-D Gramine Library OS
@@
Gramine [^2] https://gramineproject.io/ allows one to run unmodified applications inside an SGX enclave using a library OS, enabling the use of off-the-shelf software such as Python. This requires a manifest file which describes the application’s security configuration, including information on files that will be accessible, with or without integrity checking. By default, only files explicitly listed as trusted files (whose data is incorporated into the enclave hash at build time, and have their integrity checked on access), or allowed files (which can be accessed but are not subject to integrity checks) are accessible from within the enclave.
This occurs by the Gramine tooling automatically hashing the trusted files, and adding them to the enclave manifest file, which is incorporated into the SGX enclave’s MRENCLAVE. Thus, marking files as trusted prevents them from being silently replaced by a malicious host, as the subsequent hash verification will fail. This allows even security-critical files to be stored outside the enclave.
III Problem Statement
@@
We propose the term ML property cards to collectively refer to declarations like model cards and datasheets used by model providers in the interest of transparency. Our goal is to design a framework to generate verifiable ML property cards with the help of ML property attestations for training and inference. This will help attest information in ML property cards (a) for regulatory compliance, and (b) is not false or tampered. We first present the adversary model (Section III-A), desiderata for an ideal attestation mechanism (Section III-B), and limitations of current attestation mechanisms (Section III-C).
III-A Adversary Model
@@
We consider a prover ($\mathcal{P}$) that designs, trains or evaluates a model $\mathcal{M}$. During training, $\mathcal{P}$ wants to attest different properties of $\mathcal{M}$’s training to a verifier ($\mathcal{V}$). For instance, attesting that $\mathcal{M}$ was trained using $\mathcal{T}$, $\mathcal{M}_{ar}$, and $\mathcal{D}_{tr}$ (aka “proof of training”) which satisfies some required property $p$. After training, $\mathcal{P}$ evaluates $\mathcal{M}$ to check that is satisfies some minimal accuracy, fairness and robustness requirements prior to deployment. Once $\mathcal{P}$ deploys $\mathcal{M}$ as a service, clients can query the model by sending $\mathcal{I}$ via an API interface and obtain the corresponding $\mathcal{O}$. Here, $\mathcal{P}$ attests that $\mathcal{O}$ was generated $\mathcal{M}$ for a specific $\mathcal{I}$. In Section IV-A, we expand the notion of ML property cards to include a new category, inference cards to cover this type of inference-time attestation.
@@
Since $\mathcal{P}$ is untrusted, they can try to fool the attestation by trying to attest themselves despite not meeting the requirements. For instance, making false claims despite not using the expected $\mathcal{T}$ or $\mathcal{M}_{ar}$, using a different dataset than $\mathcal{D}_{tr}$ to train $\mathcal{M}$ which does not satisfy $p$, incorrect values for accuracy, fairness, and robustness, or using a different $\mathcal{O}$ for inference. We assume that $\mathcal{P}$’s entire computing platform except for the TEE is untrusted, including the privileged code such as the operating system and hypervisor are controlled by $\mathcal{P}$.
@@
We assume two roots of trust: (a) signature verification keys of external trusted organizations (e.g., CIFAR) which provides external certificates, e.g., for a test dataset, and (b) the manufacturer of the TEE (e.g., Intel) who certifies the keys used by enclaves to signi attestations. As long as $\mathcal{V}$ trusts these roots of trust, $\mathcal{V}$ can then derive trust in any attestation or external certificate that can be mapped to them.
III-B Desiderata for ML Property Attestations
We identify the following desiderata that an underlying ideal ML attestation scheme must satisfy:

### v3

@@ Abstract.
Regulations increasingly call for various assurances from machine learning (ML) model providers about their training data, training process, and model behavior. For better transparency, industry (e.g., Huggingface and Google) has adopted model cards and datasheets to describe various properties of training datasets and models. In the same vein, we introduce the notion of inference cards to describe the properties of a given inference (e.g., binding of the output to the model and its corresponding input). We coin the term ML property cards to collectively refer to these various types of cards.
@@
To prevent a malicious model provider from including false information in ML property cards, they need to be verifiable. We show how to construct verifiable ML property cards using property attestation, technical mechanisms by which a prover (e.g., a model provider) can attest to various ML properties to a verifier (e.g., an auditor). Since prior attestation mechanisms based purely on cryptography are often narrowly focused (lacking versatility) and inefficient, we need an efficient mechanism to attest different types of properties across the entire ML model pipeline.
@@
Emerging widespread support for confidential computing has made it possible to run and even train models inside hardware-assisted trusted execution environments (TEEs), which provide highly efficient attestation mechanisms. We propose Laminator, which uses TEEs to provide the first framework for verifiable ML property cards via hardware-assisted ML property attestations. Laminator is efficient in terms of overhead, scalable to large numbers of verifiers, and versatile with respect to the properties it can prove during training or inference{arxiv} [^1] ACM Conference on Data and Application Security and Privacy (CODASPY), 2025..
copyright: none
1. Introduction
@@
Machine learning (ML) models are increasingly being deployed for various high-stakes applications like medical diagnosis, bank loans, and autonomous vehicles (Christian, 2021). This has raised concerns about the various risks to data privacy, fairness and robustness (Kaur et al., 2022). Further, several emerging regulations strive to ensure that the training process and model behavior during inference are as expected (High-Level Expert Group on AI, 2019; Tabassi et al., 2019; de Miguel Beriain et al., 2022; Congress, 2022; Commission, 2024; Law, 2018; Kazim et al., 2021; House, 2020). For instance, a model provider may be required to convince a verifier (e.g., a regulator or potential customer) that a model meets the expected level of accuracy, fairness, privacy or robustness (High-Level Expert Group on AI, 2019; House, 2020; Kazim et al., 2021; Law, 2018; Commission, 2024) or distributional properties of training data comply with regulations (de Miguel Beriain et al., 2022; Congress, 2022). Privacy or confidentiality considerations may imply that model providers need to do so without disclosing the model or the training data to the verifier. The ability to demonstrate regulatory compliance without compromising confidentiality of models or training data will enable the development of a marketplace for ML models.
@@
Several companies, including Google and Huggingface, have adopted model cards (Mitchell et al., 2019) and datasheets (Gebru et al., 2021; Pushkarna et al., 2022) for better transparency of ML model and dataset properties, respectively. We propose a similar notion called inference cards for describing the properties of a specific inference, e.g., binding the output of an inference, to the input and the model. Collectively, we refer to them as ML property cards which are short documents containing information about ML properties such as intended use, risks and limitations, training procedure, datasets, and metrics for evaluation (e.g., accuracy, fairness, robustness) (Mitchell et al., 2019). These ML property cards are generated by the model provider who trains and/or deploys the model, implicitly assuming that model providers are trusted to include correct information in ML property cards. But this may not be a valid assumption, as it has been demonstrated that malicious entities can distribute models via marketplaces (Mithril-Security, 2023). Hence, ML property cards must be verifiable to ensure that their information is neither false nor tampered with.
@@
Huggingface offers a utility that allows model providers to send their model to a trusted certifier (e.g., Huggingface) to compute some metrics. The trusted certifier then creates a property card with a signature to verify that the metric is valid. However, this is not easily-extensible, being limited only to metrics that the certifier can compute. Hence, it is not possible to prove i) properties of sensitive datasets that cannot be disclosed to the certifier (e.g., demographic distribution, or certifying provenance of training dataset) ii) that a model was trained using a specific optimization (e.g., differential privacy (Abadi et al., 2016), or group fairness (Hardt et al., 2016)), or iii) that an output is the result of the model applied on a corresponding input during inference. We propose using ML property attestation, first introduced in our previous work (Duddu et al., 2024), which are technical mechanisms by which a prover (e.g., a model provider) can demonstrate relevant properties (training-time and inference-time) to a verifier (e.g., a client querying a model deployed as a service, a customer purchasing a model, or a regulatory authority).
@@
Current ML property attestation schemes have various drawbacks in terms of accuracy (Duddu et al., 2024), efficiency (Duddu et al., 2024; Garg et al., 2023), or scalability to multiple verifiers (Duddu et al., 2024). Further, prior works are limited to specific types of attestation such as proof of training (that a model was trained on some dataset) (Garg et al., 2023; Abbaszadeh et al., 2024; Shamsabadi et al., 2023; Sun and Zhang, 2023), distributional property attestations (proving distributions of attributes in training data) (Duddu et al., 2024; Chang et al., 2023), or showing that a model was trained with a specific optimization (Shamsabadi et al., 2024; Kilbertus et al., 2018; Park et al., 2022; Segal et al., 2021; Yadav et al., 2024). No existing property attestation mechanism is simultaneously efficient, scalable, and versatile.
Trusted Execution Environments (TEEs) already have the notion of hardware-assisted remote attestation to prove local system or software configuration to a remote verifier. The trusted computing research community has extended these mechanisms to more general property attestation (Sadeghi and Stüble, 2004). TEEs are widely available in commercially-available confidential computing environments (Sardar and Fetzer, 2023). Recent developments by several hardware vendors (Intel, 2024; Nvidia, 2024), have made it possible to run and train ML models efficiently inside TEEs. We take advantage of the increasing use of TEEs in ML deployments to realize efficient, scalable and versatile ML property attestation.
@@
We propose the first software framework for hardware-assisted ML property attestation. We use TEEs to establish bindings between key properties of training dataset and the model during inference and training. These bindings are signed by the TEE’s secret attestation key, generating ML property attestations. When combined with signed external certificates from trusted entities, these attestations enable verifiers to draw conclusions about models, their behavior, or their training datasets. These attestations can be then be used for verifiable ML property cards. We claim the following contributions: we
  1. (1)
@@
    propose inference cards which include properties specific to each inference (e.g., binding input-model-output) and the collective of ML property cards to include model cards, datasheets, and inference cards (Section 4)
  2. (2)
@@
    propose Laminator  [^2] Code: https://github.com/ssg-research/laminator., the first framework to generate verifiable ML property cards using hardware-assisted attestations throughout training and inference (Section 5)
  3. (3)
    using four different types of attestations, we demonstrate that Laminator is efficient, scalable, and versatile. (Section 6)
2. Background
@@
We present the ML notations used in this work (Section 2.1), details about model and dataset cards (Section 2.2), and trusted execution environments (Section 2.3).
2.1. Machine Learning Notations
@@
Consider an ML model configuration file $\mathcal{M}_{ar}$ describing the model architecture including the number and types of layers, and a training configuration file $\mathcal{T}$ which includes the objective function, learning rate, weight decay, and other hyperparameters such as number of epochs. Additionally, consider a dataset $\mathcal{D}_{tr}\sim\mathbb{D}_{\textsc{p}}$ where $\mathbb{D}_{\textsc{p}}$ is an underlying distribution satisfying some distributional properties p (for example, the ratio of females in the distribution). Since, we use i.i.d sampling to get $\mathcal{D}_{tr}$, it satisfies p. We obtain model $\mathcal{M}$ by training as per the description in $\mathcal{T}$ using $\mathcal{M}_{ar}$ on $\mathcal{D}_{tr}$, i.e., $\mathcal{M}\leftarrow\text{Train}(\mathcal{D}_{tr},\mathcal{M}_{ar},\mathcal{T})$. During inference, $\mathcal{M}$ takes an input $\mathcal{I}$ and output $\mathcal{O}$, i.e., $\mathcal{O}$leftarrow $\mathcal{M}$($\mathcal{I}$). Additionally, $\mathcal{M}$ is evaluated with respect to accuracy, fairness or robustness on some unseen test dataset $\mathcal{D}_{te}$ sampled from same distribution as $\mathcal{D}_{tr}$. We assume that both $\mathcal{D}_{tr}$ and $\mathcal{D}_{te}$ are composed of data records $\mathcal{I}$, the corresponding classification label $y$, and some sensitive attribute $z$. We summarize notations used in the rest of the paper in Table 1.
Table 1. Summary of notations and their descriptions.
@@
2.2. Model Cards and Datasheets
@@
Model cards are short documents containing additional information about different static model properties (Mitchell et al., 2019). These include properties prior to deployment of the model as a service to clients to query via an API reference. They were designed to provide better transparency into the design of ML model which includes training and evaluation of the model prior to deployment for inference. These have been extensively adopted by different companies such as Google (Google, 2024) and Huggingface (HuggingFace, 2024b). Model cards contain a variety of additional information about ML properties: i) general propertiesincluding information about the model’s intended use, risks such as bias, and other limitations ii) training propertiesincluding training datasets used, training hyperparameters, and model architecture iii) evaluation propertiesincluding information about the different test datasets and metrics used for measuring accuracy, precision, recall, fairness, and robustness.
Datasheets describe different properties corresponding to the training dataset  (Gebru et al., 2021; Pushkarna et al., 2022). Datasheets, also known as dataset cards, have been used extensively by companies such as Huggingface (HuggingFace, 2024a). They describe collection process and sources from which the raw data was collected, pre-processing including cleaning and labelling to obtain training dataset from raw data, intended use, composition of dataset, and demographic subpopulations and their distribution.
@@
2.3. Trusted Execution Environments (TEEs)
@@
TEEs (Kostiainen et al., 2010) provide an environment in which security-critical code (trusted applications, or TAs) can run without outside interference from other TAs, or the rest of the system (the Rich Execution Environment, or REE, which contains the operating system and normal software). TEEs can be described in terms of three main properties: isolation, secure storage, and remote attestation (Kostiainen et al., 2010). Isolation refers to the inability for outside software to interfere with the operation of a TA, or to read its confidential state. Secure storage refers to persistent storage made available to TAs that resists interference by other software. Remote attestation refers to the ability of a TA to prove some aspect of a TEE’s state to a remote party, e.g., that the secret key of a particular keypair is held by a TA whose code has a particular cryptographic hash.
This work uses Intel’s Software Guard Extensions (SGX) (Intel, 2014) in its evaluation, though any TEE technology that offers isolation and remote attestation can equally be used. SGX provides enclave-based TEEs, allowing applications to create a TEE (an enclave) in their own memory space and run a TA of their choice. SGX ensures that neither the host application, the OS, nor external code can access the enclave’s memory, with communication occurring through defined entry/exit points and via memory read/write operations with the host application. SGX enclaves can remotely attest themselves by obtaining a quote, which serves as a signature over TA-provided data and certifies the enclave’s identity. This identity includes MRENCLAVE (a hash of the TA’s initial memory), MRSIGNER (a hash of the public key that signed the TA), user data (a chunk of data provided by the TA), debug mode status, versioning, and other data (Intel, 2014, EREPORT).
@@
A verifier can assess the validity of a quote, then use the hash in its MRENCLAVE and knowledge of the behavior of the corresponding code to determine the TA’s state when the quote was generated. If a TA produces quotes only for specific computations, the verifier can confirm that the computation was performed and the quoted result was produced. This enables proving properties of data and computations within the TA.
2.4. Gramine Library OS
@@
Gramine [^3] https://gramineproject.io/ allows to run unmodified applications inside an SGX enclave using a library OS, enabling the use of off-the-shelf software such as Python. This requires a manifest file which describes the application’s security configuration, including information on files that will be accessible, with or without integrity checking. By default, only files explicitly listed as trusted files (whose data is incorporated into the enclave hash at build time, and have their integrity checked on access), or allowed files (which can be accessed but are not subject to integrity checks) are accessible from within the enclave. This happens when the Gramine tooling automatically hashes the trusted files and adds them to the enclave manifest, which is then incorporated into the SGX enclave’s MRENCLAVE. Thus, marking files as trusted prevents them from being replaced by a malicious host, as the hash verification will fail. This enables even security-critical files to be stored outside the enclave.
3. Problem Statement
@@
We introduce the term ML property cards to refer to transparency declarations like model cards and datasheets used by model providers. Our goal is to design a framework for generating verifiable ML property cards using ML property attestations for training and inference. This ensures the information in the cards is (a) correctly measured and (b) not falsified or tampered with. We begin by presenting the adversary model (Section 3.1), desiderata for an ideal attestation mechanism (Section 3.2), and limitations of current attestation mechanisms (Section 3.3).
3.1. Adversary Model
@@
We consider a prover $\mathcal{P}$ that trains, evaluates, or deploys a model $\mathcal{M}$. We assume $\mathcal{P}$ has their own train and test dataset indicated as $\mathcal{D}^{prov}_{tr}$ and $\mathcal{D}^{prov}_{te}$. During training, $\mathcal{P}$ wants to attest different properties of $\mathcal{M}$’s training to a verifier $\mathcal{V}$. For instance, attesting that $\mathcal{M}$ was trained using $\mathcal{T}$, $\mathcal{M}_{ar}$, and $\mathcal{D}_{tr}$ (a.k.a. “proof of training”). After training, $\mathcal{P}$ evaluates $\mathcal{M}$ to check that it satisfies some minimal accuracy, fairness and robustness requirements prior to deployment. Once $\mathcal{P}$ deploys $\mathcal{M}$ as a service, clients can query $\mathcal{M}$ by sending input $\mathcal{I}$ via an API interface and obtain the corresponding output $\mathcal{O}$. Here, $\mathcal{P}$ attests that $\mathcal{O}$ was generated by $\mathcal{M}$ for a specific $\mathcal{I}$. In Section 4.1, we expand the notion of ML property cards to include a new category, inference cards to cover this type of inference-time attestation.
@@
$\mathcal{V}$ may not trust $\mathcal{P}$. If $\mathcal{P}$ can subvert the attestation mechanism, it can fool $\mathcal{V}$ into believing that some property holds when in fact it does not. For instance, $\mathcal{P}$ can make false claims despite not using the expected $\mathcal{T}$ or $\mathcal{M}_{ar}$, using a different dataset than $\mathcal{D}_{tr}$ which does not satisfy $p$, incorrect values for accuracy, fairness, and robustness, or using a different $\mathcal{O}$ for inference. We assume that $\mathcal{P}$’s entire computing platform except for the TEE is untrusted, including the privileged code such as the operating system and hypervisor, which are controlled by $\mathcal{P}$.
@@
We assume two roots of trust, signature verification keys belonging to (a) the manufacturer of the TEE (e.g., Intel) who certifies the keys used by enclaves to sign attestations, and (b) trusted certifiers (e.g., CIFAR) that provide attestations of computational properties (with trust in the property grounded in the trusted certifier rather than the TEE), and certificates of non-computational properties, e.g., data provenance. As long as $\mathcal{V}$ trusts them, $\mathcal{V}$ can trust any attestation or certificates linked to them.
3.2. Desiderata for ML Property Attestations
We identify the following desiderata for ML property attestation: